OpenStudy (anonymous):

In what interval is the function f(x) = (square root)(x^2+5x+4(end square root)

OpenStudy (debbieg):

in what interval is \(\large f(x)=\sqrt{x^2+5x+4}\)...... what?

OpenStudy (anonymous):

what's the interval notation that it's in

OpenStudy (anonymous):

an explanation would be nice too :)

OpenStudy (debbieg):

I don't know what you mean, "in what interval". Usually we would ask something like: "in what interval is f(x) increasing?" ... or... "in what interval is f(x) decreasing?" Or maybe you mean the domain of f(x)?? I don't know what the question means, "in what interval is f(x)."

OpenStudy (debbieg):

And I only give explanations, not answers. :) but I can't explain without knowing what the question is asking. Have you given the EXACT question, as it is asked?

OpenStudy (anonymous):

i mean what's the interval notation in regards to domain, and that is the exact question, i was confused at first too

OpenStudy (debbieg):

Wow that goofy wording to be asking for domain, lol... but ok, THAT we can work with. OK, so you have some stuff under a square root, right? So in terms of domain, what do you have to worry about? What condition has to be true?

OpenStudy (anonymous):

i have no clue...

OpenStudy (debbieg):

If \(\large f(x)=\sqrt{x}\), what is the domain? What must be true about x?

OpenStudy (anonymous):

[0,infinity)

OpenStudy (debbieg):

OK, right... must have \(x \ge0\). good. so here, we have the function: \(\large f(x)=\sqrt{x^2+5x+4}\) so the stuff under the square root HAS to be non-negative, that's what you need to make sure of. In other words, when you figure out for what x's it is true that: \(\large x^2+5x+4 \ge0\) ..then you'll have your domain. Follow so far?

OpenStudy (anonymous):

I didn't think about it that way, thanks so much

OpenStudy (debbieg):

So the x-values where \(\large x^2+5x+4 \ge0\) is exactly what your domain is. Do you have any thoughts on how to find where that is true? How do you solve an inequality like that, for a quadratic?

OpenStudy (debbieg):

ok, are you good now?

OpenStudy (anonymous):

\[\frac{ x-7 }{ x ^{2}-1 }\]

OpenStudy (anonymous):

Can you help me determine the domain and range of that one? It also asks for horizontal and vertical asymptotes

OpenStudy (debbieg):

Sure.... domain first. It's a rational function, so what do we have to worry about on THIS one? (just like for a square root, we have to worry about one particular thing, when we have a rational function, there is one specific condition that we need to worry about when it comes to domain...)

OpenStudy (anonymous):

The denominator cannot be 0, or there'd be a divide by zero error?

OpenStudy (debbieg):

Exactly. So, find where the den'r=0, and you know what's NOT in your domain. That's the only domain "issue" with this kind of function, so everything else IS in the domain. In other words, the domain is all real numbers EXCEPT those that make the den'r=0.

OpenStudy (debbieg):

So you just need to solve \(x^2-1=0\) The solutions to that are NOT allowed in the domain.

OpenStudy (anonymous):

So 1 & -1?

OpenStudy (debbieg):

Exactly! which leads us to.... the vertical asymptotes. Because those occur exactly at the x-values that make the den'r = 0.

OpenStudy (anonymous):

x=1, x=-1 for vertical asymptotes?

OpenStudy (debbieg):

Yup, exactly.

OpenStudy (debbieg):

For horizontal asymptotes, there are some easy rules that depend on the degree of the num'r vs. the degree of the den'r. Here, the degr of the num < degr of the den'r.... in that case, you will always get a HA at y=0 (the x-axis). Think about it, and it will make sense: as x gets BIG, the den'r will "take over" the num'r, because den'r has higher degree. So it bullies the num'r out of the way :) and the ratio goes to 0.

OpenStudy (anonymous):

is that the only horizontal asymptote?

OpenStudy (debbieg):

There is NEVER more than one HA. :) (And there might not be any.) Why? Because the HA is the END behavior of the function.... what y-value the function is "settling in on" for very large x. Well, at some point, it has to "settle"! E.g., it can't go to two DIFFERENT y-values, right? (If you understand what makes a relation a function, then you know why - if it settled at two y-values, then it wouldn't be a function).

OpenStudy (anonymous):

yeah, i get it, thanks so much

OpenStudy (debbieg):

You're welcome, happy to help. :)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."